Which of the ordered pairs is a solution for the equation 2x – 3y = 6?
A) (0, 2)
B) (–3, 0)
C) (–3, –2)
D) (0, –2)

Answer by funmath(2873)

Which of the ordered pairs is a solution for the equation 2x – 3y = 6?
A) (x,y)=(0, 2)
Substitute 0 in for x and 2 in for y, simplify and see if both sides of the equation equal. If they do, it is a solution. If not, it's not.
2(0)-3(2)=6
0-6=6
-6=6 Not true, so (0,2) is not a solution.
:
B) (–3, 0)
2(-3)-3(0)=6
-6-0=6
-6=6 Not true, so (-3,0) is not a solution.
:
C) (–3, –2)
2(-3)-3(-2)=6
-6+6=6
0=6 Not true, so (-3,-2) is not a solution.
:
D) (0, –2)
2(0)-3(-2)=6
0+6=6
6=6 True, so (0,-2) is a solution.
:
There are an infinite number of solutions, so you have to check each set of coordinates.
Happy Calculating!!!!